OpenStudy (qqstory):

Find the domain of each function. f(x) = -5x+4

OpenStudy (anonymous):

all real numbers

OpenStudy (anonymous):

d= all real numbers

OpenStudy (anonymous):

Domain is R

OpenStudy (qqstory):

{xl x e r}?

OpenStudy (anonymous):

domain is r it should be R(R denotes the set of all real numbers)

OpenStudy (anonymous):

{X: whaer x

OpenStudy (qqstory):

uh i want a complete answer lol..

OpenStudy (anonymous):

Domain is simple here.. For every real value of x you will get f(x) a defined value.. Meaning f(x) is defined for every real value of x.. Real value means value taken from the set of Real Numbers.. So here Domain is all reals.. you can show it by: \[Domain : \mathbb{R}\]

OpenStudy (qqstory):

my teacher wants me to write it something like this {X l X E R} is this correct?

OpenStudy (anonymous):

Yes, this is also correct..

OpenStudy (anonymous):

\[Domain : \{x : x \in \mathbb{R}\}\]

OpenStudy (anonymous):

he's talking about the set notation, and then the graph

OpenStudy (anonymous):

something like that @waterineyes

OpenStudy (anonymous):

No: he asked to find the domain and write it in particular set notation.. What set notation is this @panlac01 ?/ I always have confusion in what is roaster form and what is set builder form.. Can you tell??

OpenStudy (anonymous):

http://www.purplemath.com/modules/setnotn.htm something students here in the US learn under set theory.

OpenStudy (qqstory):

how to solve this one? f(x)=x/x^2+1

OpenStudy (qqstory):

is this correct? {xlx\[\neq\]0}

OpenStudy (anonymous):

you must meet a condition for a function @qqstory that is denominator cannot be zero. you can solve that by x^2+1=0

OpenStudy (anonymous):

and for numerator, it cannot have a square root

OpenStudy (qqstory):

o_O so am i right? {xl x \[\neq\] 0}